OpenStudy (kitkat1):

Medal for best answer!!! Select the polynomial that is a perfect square trinomial. 9x2 + 9x + 1 36b2 − 24b + 8 16x2 + 24x + 9 4a2 − 10a + 25

OpenStudy (welshfella):

try factoring them

OpenStudy (kitkat1):

It's either 1st or 3rd

OpenStudy (kitkat1):

Both 1st and 3rd end up to be the middle number

OpenStudy (kitkat1):

@tkhunny

OpenStudy (welshfella):

Hint the first and last terms will be perfect squares - and the middle number must fit in as you said

OpenStudy (welshfella):

you are right its either 1 or 3

OpenStudy (kitkat1):

They both qualify I don't know which one is better

OpenStudy (welshfella):

no if you try factorinbg them you'[ll see which is the right one

OpenStudy (kitkat1):

Idk how to factor

OpenStudy (welshfella):

9x^2 ~ 9x + 1 sqrt 9x^2 = 3x and sqrt 1 = 1 so lets write (3x + 1)(3x + 1) does that expand to 9x^2 + 9x + 1?

OpenStudy (kitkat1):

Yes

OpenStudy (welshfella):

(3x + 1)(3x + 1) = 3x(3x + 1) + 1(3x + 1) can you continue?

OpenStudy (welshfella):

= 9x^2 + 3x + 3x + 1

OpenStudy (welshfella):

- it does not... ]'the middle term is 6x

OpenStudy (kitkat1):

Okay so the answer is the 1st one?

OpenStudy (welshfella):

NO!! because (3x + 1)^2 is not = first one - its not a perfect square

OpenStudy (kitkat1):

3rd one?

OpenStudy (welshfella):

yes 9X62 + 6x + 1 is a perfect square but 9x^2 + 9x + 1 is

OpenStudy (welshfella):

do you understand the reasoning?

OpenStudy (kitkat1):

Yes I have one more that I am not sure can you check it?

OpenStudy (kitkat1):

Given the following perfect square trinomial, fill in the missing term. x2 − 16x +

OpenStudy (welshfella):

please post it separately

OpenStudy (kitkat1):

Okay I will tag you